OpenStudy (anonymous):

A recipe that makes 3 servings calls for 3/5 cup of apple juice and 3/4 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (anonymous):

A. 4 1/20 B. 1 7/20 C. 2/3 D. 9/20

OpenStudy (anonymous):

35+34 we need first find a common denominator..a number both 5 and 4 go into evenly..20 seems about right so we need to multiply the whole first fraction by 4 to make the denominator 20...and the whole second fraction by 5 to make that denominator 20 so 44×35+55×34 this will come out to 1220+1520 which will equal 2720

OpenStudy (anonymous):

B or d i thiank

OpenStudy (texaschic101):

3/5 + 3/4 = 12/20 + 15/20 = 27/20 (total juice) 3/(27/20) = 1/ (x) --- 3 servings to 27/20 juice = 1 serving to x juice cross multiply (3)(x) = (27/20)(1) 3x = 27/20 -- divide by 3 x = (27/20) / 3 x = 27/20 * 1/3 x = 27/60 which reduces to 9/20